Two grams of hydrogen diffuse from a container in 10 minutes how many grams of oxygen would diffuse through the same container in the same time under similar conditions

Grade:6

1 Answers

Aashna
6 years ago
Rate of diffusion is inversally proportionAl to square root of mass. R1/R2=√M2/√M1R1 = 10R2 = ?M1 = 2(hydrogen gas)M2 = 32(oxygen gas)R2 = 2.5mins